In a linear series plate modulated amplifier, the plate supply vottage is 300V and the D.C plate current under unmodulated condition is 20A, the sinusoidal modulating voltage which appears in the plate current of modulated complifier has amplitude of 150V and the unmodulated output Carrier power is 4.5KW, find; (i) modulation index (ii) plate current efficiency (iii) Carrier power under modulated condition (iv) plate dissipation under unmodulated and modulated conditions
1
Expert's answer
2019-11-06T08:05:59-0500

(i) Modulation index:


"m=\frac{M}{A}=\frac{150}{300}=0.5."

(ii) Plate current efficiency. The D.C plate current under unmodulated condition is 20A, the unmodulated output carrier power is 4.5KW, therefore the current is


"I=\frac{4500}{300}=15\text{ A},"

and the efficiency


"a=\frac{15}{20}=0.75."

(iii) Carrier power under modulated condition is


"P_m=300\cdot20=6000\text{ W}."

(iv) Plate dissipation under unmodulated and modulated conditions:


"D_m=4500(1-0.75)=1125\text{ W},\\
D_u=6000(1-0.75)=1500\text{ W}."